A new proof of the Scott correction

Wolfgang Spitzer UC Davis, Mathematics

We introduce new coherent states which we use to prove local semi-classical results for eigenvalues of Schrödinger operators. From there we prove the Scott correction for molecules. This is a joint work with Jan Philip Solovej.